# Latent Concept-based Explanation of NLP Models

### v2

@@ H.1 Experimental Setup
H.2 Qualitative Evaluation
H.2.1 Correct prediction with correct gold label
@@
H.2.2 Wrong prediction with correct gold label
@@
H.3 Module Specific Evaluation
H.3.1 ConceptDiscoverer
@@ H.3.3 ConceptMapper
Appendix I NLI Task
I.1 Experimental Setup
@@ I.2 Qualitative Evaluation
I.3 Module Specific Evaluation
I.3.1 ConceptDiscoverer
@@ I.3.3 ConceptMapper
Appendix J LLama2
J.1 Experimental Setup
@@
J.2 Sentiment Classification Task
J.2.1 ConceptDiscoverer
Compared to the BERT, RoBERTa, and XLMR models (Table 7), the Llama2 model has fewer class-based clusters at the last layer(See Table 21). There are around 67% class-based clusters detected at the last layer for the Llama2 model. The BERT, RoBERTa, and XLMR models have 78%, 95%, and 94% class-based clusters at the last layer.
@@
J.3 Toxicity Classification Task
J.3.1 ConceptDiscoverer

### v3

@@ F.1 Experimental Setup
F.2 Qualitative Evaluation
F.2.1 Correct prediction with correct gold label
@@
F.2.2 Wrong prediction with correct gold label
@@
F.3 Module Specific Evaluation
F.3.1 ConceptDiscoverer
@@ F.3.3 ConceptMapper
Appendix G NLI Task
G.1 Experimental Setup
@@ G.2 Qualitative Evaluation
G.3 Module Specific Evaluation
G.3.1 ConceptDiscoverer
@@ G.3.3 ConceptMapper
Table 20: Top 1, 2, and 5 accuracy of  ConceptMapper in mapping a representation to the correct latent concept for the MNLI task. The top-5 performance reaches above 90% for all models demonstrating that the correct latent concept is among the top probable latent concepts of ConceptMapper.
Appendix H LLama2
H.1 Experimental Setup
@@
H.2 Sentiment Classification Task
H.2.1 ConceptDiscoverer
Compared to the BERT, RoBERTa, and XLMR models (Table 8), the Llama2 model has fewer class-based clusters at the last layer(See Table 21). There are around 67% class-based clusters detected at the last layer for the Llama2 model.
@@ H.2.3 ConceptMapper
H.3 Toxicity Classification Task
H.3.1 ConceptDiscoverer